# Survival Analysis with Adversarial Regularization

### v3

@@ VI-D2 Population Survival Curve
We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations (Tables III and II). We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes. More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case perturbation for each dataset are available in the code repository linked in the abstract.
@@ VII Conclusions
This paper introduces CROWN-IBP regularization for enhancing adversarial robustness in fully-parametric SA using NN models. We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods such as Gaussian noise, FGSM, and PGD, as well as the novel worst-case perturbation.
Our empirical results show that CROWN-IBP regularization consistently improves performance in SA across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1. Compared to conventional adversarial regularization techniques, SAWAR demonstrates superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1% to 150% across all $\epsilon$-perturbation magnitudes.
In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration. This approach effectively mitigates data uncertainty in SA and improves generalization across unseen datasets.
Future work will include the use the Weibull CPH model to enable time varying hazard rates. This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
@@ Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
%$\%\Delta$ 29.46 48.18 56.27 72.23 102.44 133.35 152.17 140.33 91.78 21.17 9.63 1.07
TABLE IV: The relative percent change in Concordance Index metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets. A higher relative percent change is better.
@@
Aids2 DRAFT 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
Noise 0.492 0.494 0.496 0.5 0.5 0.498 0.507 0.526 0.556 0.556 0.554 0.554
FGSM 0.498 0.495 0.498 0.497 0.5 0.502 0.508 0.529 0.554 0.557 0.555 0.553
PGD 0.498 0.497 0.497 0.5 0.501 0.501 0.507 0.531 0.557 0.558 0.556 0.554
SAWAR 0.567 0.575 0.577 0.579 0.58 0.58 0.579 0.579 0.579 0.579 0.579 0.579
Framingham DRAFT 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
Noise 0.606 0.616 0.626 0.635 0.642 0.649 0.659 0.674 0.683 0.685 0.685 0.685
FGSM 0.607 0.611 0.617 0.622 0.626 0.635 0.652 0.669 0.682 0.686 0.687 0.686
PGD 0.567 0.578 0.588 0.599 0.61 0.624 0.642 0.659 0.68 0.688 0.688 0.687
SAWAR 0.694 0.697 0.699 0.701 0.702 0.704 0.704 0.705 0.705 0.704 0.704 0.704
LeukSurv DRAFT 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
Noise 0.552 0.557 0.562 0.566 0.569 0.574 0.587 0.61 0.628 0.627 0.621 0.617
FGSM 0.546 0.549 0.554 0.563 0.561 0.566 0.583 0.602 0.624 0.624 0.621 0.62
PGD 0.537 0.536 0.538 0.549 0.546 0.558 0.577 0.598 0.62 0.622 0.619 0.617
SAWAR 0.483 0.498 0.51 0.532 0.553 0.583 0.589 0.607 0.617 0.623 0.632 0.636
TRACE DRAFT 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
Noise 0.565 0.588 0.613 0.64 0.668 0.692 0.716 0.73 0.733 0.735 0.734 0.735
FGSM 0.553 0.577 0.603 0.63 0.658 0.688 0.713 0.722 0.727 0.73 0.733 0.734
PGD 0.557 0.582 0.609 0.636 0.664 0.696 0.723 0.731 0.733 0.732 0.733 0.733
SAWAR 0.73 0.733 0.736 0.738 0.739 0.739 0.74 0.74 0.739 0.738 0.738 0.737
dataDIVAT1 DRAFT 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
Noise 0.534 0.539 0.551 0.571 0.587 0.612 0.636 0.656 0.654 0.64 0.635 0.631
FGSM 0.575 0.585 0.594 0.606 0.617 0.628 0.638 0.644 0.651 0.643 0.638 0.635
PGD 0.517 0.529 0.548 0.568 0.591 0.618 0.636 0.654 0.657 0.651 0.647 0.644
SAWAR 0.627 0.642 0.652 0.658 0.659 0.659 0.659 0.658 0.657 0.656 0.656 0.655
flchain DRAFT 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
Noise 0.096 0.099 0.104 0.112 0.127 0.164 0.335 0.768 0.912 0.921 0.923 0.923
FGSM 0.102 0.116 0.159 0.288 0.599 0.852 0.91 0.917 0.923 0.927 0.927 0.925
PGD 0.102 0.127 0.235 0.538 0.824 0.904 0.917 0.921 0.925 0.929 0.929 0.929
SAWAR 0.84 0.925 0.927 0.929 0.93 0.93 0.929 0.929 0.93 0.93 0.931 0.931
prostate DRAFT 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
Noise 0.49 0.503 0.514 0.524 0.533 0.541 0.553 0.566 0.576 0.572 0.577 0.576
FGSM 0.514 0.522 0.526 0.536 0.541 0.549 0.558 0.573 0.582 0.596 0.596 0.596
PGD 0.522 0.526 0.532 0.535 0.54 0.548 0.557 0.567 0.576 0.593 0.591 0.593
SAWAR 0.575 0.584 0.598 0.6 0.608 0.632 0.638 0.638 0.639 0.641 0.642 0.644
retinopathy DRAFT 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
Noise 0.526 0.545 0.561 0.57 0.581 0.588 0.588 0.6 0.603 0.6 0.604 0.606
FGSM 0.47 0.485 0.51 0.536 0.569 0.578 0.591 0.6 0.599 0.602 0.606 0.607
PGD 0.471 0.481 0.509 0.53 0.567 0.577 0.591 0.599 0.598 0.601 0.606 0.607
SAWAR 0.588 0.597 0.611 0.616 0.619 0.615 0.621 0.62 0.62 0.619 0.617 0.617
stagec DRAFT 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
Noise 0.471 0.481 0.505 0.49 0.514 0.534 0.582 0.697 0.721 0.712 0.688 0.697
FGSM 0.423 0.438 0.447 0.471 0.51 0.567 0.62 0.663 0.688 0.678 0.688 0.688
PGD 0.462 0.486 0.471 0.495 0.51 0.567 0.611 0.668 0.692 0.688 0.697 0.697
SAWAR 0.428 0.447 0.481 0.514 0.567 0.639 0.673 0.692 0.702 0.692 0.692 0.707
zinc DRAFT 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
Noise 0.315 0.329 0.339 0.376 0.424 0.545 0.704 0.778 0.791 0.805 0.806 0.81
FGSM 0.339 0.357 0.384 0.439 0.55 0.657 0.738 0.776 0.787 0.804 0.809 0.812
PGD 0.333 0.351 0.384 0.439 0.55 0.671 0.738 0.776 0.79 0.8 0.809 0.808
SAWAR 0.383 0.464 0.533 0.649 0.707 0.746 0.766 0.779 0.789 0.787 0.795 0.8
@@
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
%$\%\Delta$44.13 -45.46 -46.68 -47.57 -47.86 -47.10 -45.03 -40.89 -33.06 -20.29 -11.34 -1.68
TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets. A lower relative percent change is better.
@@
Aids2 DRAFT 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
Noise 0.254 0.253 0.25 0.245 0.236 0.22 0.196 0.161 0.13 0.121 0.121 0.122
FGSM 0.252 0.249 0.244 0.236 0.223 0.203 0.176 0.145 0.125 0.121 0.121 0.121
PGD 0.252 0.248 0.243 0.234 0.221 0.2 0.172 0.143 0.124 0.121 0.121 0.121
SAWAR 0.126 0.124 0.124 0.124 0.123 0.123 0.123 0.123 0.123 0.123 0.123 0.124
Framingham DRAFT 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
Noise 0.818 0.818 0.816 0.809 0.778 0.681 0.495 0.294 0.174 0.128 0.119 0.116
FGSM 0.818 0.816 0.81 0.791 0.728 0.58 0.369 0.208 0.14 0.12 0.116 0.115
PGD 0.818 0.818 0.817 0.811 0.784 0.683 0.461 0.235 0.138 0.117 0.115 0.115
SAWAR 0.165 0.152 0.141 0.133 0.127 0.122 0.118 0.116 0.114 0.114 0.113 0.113
LeukSurv DRAFT 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
Noise 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.157 0.134 0.12 0.114
FGSM 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.161 0.148 0.122 0.114 0.111
PGD 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.16 0.145 0.12 0.114 0.111
SAWAR 0.154 0.15 0.144 0.136 0.128 0.12 0.115 0.112 0.11 0.107 0.107 0.106
TRACE DRAFT 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
Noise 0.634 0.634 0.632 0.622 0.596 0.536 0.435 0.319 0.23 0.183 0.172 0.167
FGSM 0.634 0.634 0.634 0.631 0.613 0.555 0.434 0.294 0.212 0.177 0.171 0.168
PGD 0.634 0.634 0.632 0.621 0.58 0.491 0.362 0.245 0.189 0.171 0.167 0.166
SAWAR 0.282 0.254 0.231 0.213 0.198 0.186 0.177 0.17 0.166 0.163 0.163 0.163
dataDIVAT1 DRAFT 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
Noise 0.749 0.749 0.748 0.748 0.743 0.72 0.626 0.413 0.248 0.198 0.192 0.192
FGSM 0.749 0.749 0.748 0.747 0.74 0.699 0.547 0.322 0.215 0.193 0.192 0.192
PGD 0.749 0.749 0.749 0.748 0.746 0.724 0.582 0.305 0.192 0.184 0.186 0.187
SAWAR 0.266 0.241 0.224 0.212 0.202 0.193 0.187 0.183 0.18 0.178 0.178 0.178
flchain DRAFT 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
Noise 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.807 0.25 0.087 0.063 0.052
FGSM 0.831 0.831 0.831 0.83 0.83 0.828 0.707 0.146 0.094 0.06 0.054 0.051
PGD 0.831 0.831 0.831 0.83 0.83 0.797 0.298 0.114 0.079 0.055 0.051 0.05
SAWAR 0.522 0.161 0.076 0.063 0.057 0.054 0.052 0.051 0.05 0.05 0.049 0.049
prostate DRAFT 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
Noise 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.477 0.371 0.325 0.296
FGSM 0.53 0.53 0.53 0.53 0.53 0.53 0.529 0.509 0.413 0.29 0.257 0.246
PGD 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.502 0.396 0.28 0.254 0.244
SAWAR 0.508 0.482 0.441 0.393 0.348 0.309 0.278 0.251 0.229 0.214 0.209 0.205
retinopathy DRAFT 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
Noise 0.662 0.657 0.645 0.622 0.583 0.523 0.432 0.326 0.245 0.204 0.196 0.195
FGSM 0.656 0.646 0.626 0.597 0.556 0.493 0.401 0.301 0.232 0.201 0.196 0.197
PGD 0.656 0.645 0.625 0.596 0.554 0.49 0.397 0.299 0.231 0.201 0.197 0.198
SAWAR 0.526 0.484 0.436 0.385 0.334 0.29 0.254 0.228 0.211 0.203 0.202 0.202
stagec DRAFT 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
Noise 0.505 0.505 0.505 0.504 0.502 0.485 0.432 0.338 0.216 0.175 0.181 0.192
FGSM 0.505 0.505 0.505 0.503 0.494 0.458 0.382 0.268 0.174 0.161 0.169 0.178
PGD 0.505 0.505 0.505 0.503 0.493 0.454 0.377 0.261 0.169 0.159 0.167 0.176
SAWAR 0.491 0.476 0.45 0.408 0.354 0.296 0.24 0.195 0.168 0.16 0.162 0.167
zinc DRAFT 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
Noise 0.908 0.907 0.907 0.904 0.893 0.838 0.625 0.32 0.168 0.109 0.096 0.089
FGSM 0.907 0.907 0.904 0.892 0.839 0.659 0.365 0.194 0.124 0.095 0.089 0.086
PGD 0.907 0.907 0.903 0.89 0.832 0.64 0.348 0.187 0.122 0.094 0.088 0.087
SAWAR 0.72 0.609 0.476 0.352 0.26 0.198 0.159 0.132 0.113 0.1 0.096 0.094

### v4

@@ VII Conclusions
This paper introduces a robust training objective based on CROWN-IBP regularization to enhance adversarial robustness in fully-parametric NN-based SA models, improving both model calibration and generalization. We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods like Gaussian noise, FGSM, and PGD, as well as SOTA models such as DRAFT and AAE-Cox.
Our empirical results demonstrate that SAWAR consistently improves SA performance across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1. Compared to traditional adversarial regularization techniques, SAWAR shows superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1
In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration, effectively mitigating data uncertainty in SA and improving generalization on unseen datasets.
Future work will involve extending the approach to the Weibull CPH model to enable time-varying hazard rates, which will require adapting auto_LiRPA to handle power operations such as $t^{k}$, where $k$ is a learnable parameter.
@@ Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
%$\%\Delta$ 72.8 79.69 92.54 90.82 81.56 66.21 41.01 12.83 3.19 1.88 1.72 1.6
TABLE IV: The relative percent change in the Concordance Index metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the worst-case adversarial attack. A higher relative percent change is better.
@@
Dataset Algorithm
Aids2 DRAFT 0.516 0.519 0.522 0.525 0.527 0.53 0.536 0.551 0.566 0.572 0.572 0.57
Noise 0.513 0.515 0.518 0.521 0.525 0.529 0.535 0.548 0.568 0.571 0.57 0.569
FGSM 0.512 0.516 0.52 0.523 0.526 0.532 0.537 0.55 0.565 0.569 0.568 0.566
PGD 0.506 0.509 0.513 0.518 0.523 0.528 0.535 0.552 0.567 0.569 0.568 0.567
AAE-Cox 0.504 0.508 0.507 0.506 0.505 0.501 0.499 0.504 0.51 0.561 0.573 0.573
SAWAR 0.565 0.568 0.568 0.569 0.569 0.57 0.57 0.57 0.57 0.57 0.569 0.569
Framingham DRAFT 0.618 0.634 0.649 0.663 0.676 0.688 0.699 0.708 0.716 0.72 0.721 0.722
Noise 0.606 0.622 0.637 0.653 0.666 0.679 0.691 0.703 0.713 0.718 0.719 0.719
FGSM 0.582 0.598 0.611 0.626 0.641 0.655 0.671 0.691 0.708 0.715 0.715 0.715
PGD 0.539 0.559 0.582 0.605 0.627 0.647 0.669 0.693 0.71 0.717 0.717 0.717
AAE-Cox 0.514 0.522 0.529 0.533 0.539 0.553 0.569 0.57 0.582 0.73 0.733 0.733
SAWAR 0.721 0.725 0.729 0.731 0.733 0.734 0.735 0.736 0.736 0.737 0.737 0.737
LeukSurv DRAFT 0.589 0.593 0.596 0.598 0.601 0.602 0.608 0.62 0.63 0.632 0.633 0.633
Noise 0.545 0.548 0.545 0.547 0.551 0.56 0.581 0.609 0.63 0.628 0.626 0.624
FGSM 0.511 0.514 0.52 0.525 0.535 0.549 0.579 0.612 0.637 0.638 0.636 0.635
PGD 0.498 0.498 0.501 0.51 0.521 0.543 0.576 0.616 0.64 0.643 0.64 0.638
AAE-Cox 0.559 0.552 0.542 0.555 0.548 0.55 0.548 0.537 0.545 0.631 0.658 0.656
SAWAR 0.495 0.51 0.524 0.536 0.552 0.57 0.588 0.609 0.633 0.658 0.669 0.673
TRACE DRAFT 0.581 0.605 0.633 0.66 0.685 0.708 0.726 0.736 0.741 0.743 0.744 0.745
Noise 0.576 0.598 0.621 0.645 0.669 0.691 0.712 0.728 0.737 0.742 0.744 0.745
FGSM 0.581 0.603 0.629 0.653 0.674 0.697 0.717 0.732 0.739 0.743 0.744 0.744
PGD 0.571 0.595 0.621 0.646 0.668 0.691 0.714 0.73 0.739 0.743 0.744 0.745
AAE-Cox 0.432 0.44 0.447 0.451 0.46 0.47 0.489 0.524 0.628 0.743 0.746 0.747
SAWAR 0.714 0.722 0.728 0.733 0.735 0.737 0.739 0.742 0.744 0.746 0.747 0.747
dataDIVAT1 DRAFT 0.573 0.585 0.598 0.611 0.625 0.636 0.646 0.653 0.657 0.657 0.656 0.655
Noise 0.532 0.541 0.552 0.562 0.575 0.589 0.603 0.619 0.631 0.636 0.635 0.635
FGSM 0.546 0.555 0.567 0.578 0.588 0.6 0.612 0.626 0.641 0.647 0.646 0.646
PGD 0.523 0.534 0.545 0.557 0.569 0.581 0.599 0.621 0.641 0.648 0.648 0.648
AAE-Cox 0.597 0.589 0.589 0.585 0.58 0.571 0.571 0.566 0.596 0.662 0.663 0.663
SAWAR 0.64 0.645 0.649 0.654 0.658 0.661 0.664 0.665 0.667 0.668 0.669 0.67
flchain DRAFT 0.109 0.111 0.115 0.122 0.133 0.158 0.239 0.566 0.905 0.917 0.92 0.921
Noise 0.166 0.131 0.111 0.123 0.149 0.224 0.433 0.792 0.911 0.917 0.919 0.918
FGSM 0.115 0.153 0.245 0.431 0.72 0.885 0.91 0.914 0.918 0.922 0.922 0.922
PGD 0.165 0.268 0.457 0.744 0.876 0.904 0.912 0.917 0.92 0.923 0.923 0.923
AAE-Cox 0.527 0.648 0.668 0.553 0.394 0.511 0.51 0.476 0.11 0.179 0.925 0.926
SAWAR 0.593 0.684 0.866 0.918 0.922 0.925 0.926 0.927 0.927 0.927 0.927 0.927
prostate DRAFT 0.402 0.416 0.428 0.444 0.469 0.505 0.542 0.585 0.627 0.653 0.661 0.668
Noise 0.433 0.44 0.448 0.451 0.463 0.48 0.492 0.511 0.537 0.558 0.561 0.562
FGSM 0.448 0.453 0.46 0.465 0.475 0.486 0.502 0.515 0.541 0.564 0.569 0.571
PGD 0.447 0.455 0.459 0.467 0.474 0.487 0.509 0.53 0.564 0.585 0.588 0.588
AAE-Cox 0.41 0.411 0.406 0.41 0.407 0.403 0.399 0.391 0.414 0.646 0.686 0.691
SAWAR 0.608 0.623 0.637 0.652 0.665 0.671 0.672 0.673 0.667 0.663 0.66 0.657
retinopathy DRAFT 0.553 0.568 0.578 0.597 0.616 0.632 0.645 0.653 0.663 0.666 0.667 0.669
Noise 0.573 0.591 0.605 0.62 0.632 0.644 0.653 0.661 0.666 0.667 0.668 0.668
FGSM 0.575 0.592 0.604 0.615 0.625 0.633 0.642 0.647 0.651 0.656 0.657 0.659
PGD 0.571 0.589 0.599 0.61 0.621 0.63 0.64 0.647 0.651 0.655 0.656 0.657
AAE-Cox 0.494 0.495 0.506 0.504 0.514 0.564 0.577 0.592 0.62 0.657 0.652 0.648
SAWAR 0.668 0.669 0.67 0.666 0.662 0.66 0.656 0.654 0.653 0.65 0.648 0.647
stagec DRAFT 0.358 0.378 0.382 0.406 0.425 0.449 0.454 0.475 0.489 0.504 0.507 0.512
Noise 0.353 0.373 0.39 0.407 0.436 0.466 0.485 0.5 0.53 0.544 0.549 0.555
FGSM 0.329 0.346 0.368 0.389 0.416 0.429 0.443 0.471 0.488 0.502 0.503 0.51
PGD 0.341 0.352 0.381 0.399 0.419 0.431 0.442 0.47 0.49 0.496 0.498 0.505
AAE-Cox 0.393 0.397 0.397 0.411 0.417 0.409 0.405 0.429 0.469 0.523 0.543 0.54
SAWAR 0.393 0.401 0.413 0.425 0.436 0.461 0.488 0.506 0.534 0.543 0.548 0.558
zinc DRAFT 0.262 0.27 0.284 0.296 0.317 0.355 0.44 0.579 0.712 0.755 0.765 0.77
Noise 0.318 0.328 0.343 0.366 0.401 0.466 0.562 0.661 0.734 0.766 0.773 0.776
FGSM 0.377 0.403 0.439 0.49 0.557 0.626 0.685 0.737 0.762 0.78 0.782 0.783
PGD 0.384 0.406 0.443 0.495 0.559 0.632 0.695 0.741 0.77 0.778 0.78 0.781
AAE-Cox 0.226 0.231 0.228 0.236 0.234 0.249 0.265 0.306 0.445 0.724 0.754 0.759
SAWAR 0.667 0.725 0.763 0.785 0.785 0.785 0.781 0.778 0.778 0.776 0.776 0.774
@@
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
%$\%\Delta$37.55 -42.51 -48.08 -53.15 -55.48 -55.34 -51.51 -42.94 -28.8 -10.87 -4.78 -0.65
TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the worst-case adversarial attack. A lower relative percent change is better.
@@
Dataset Algorithm
Aids2 DRAFT 0.265 0.262 0.258 0.252 0.242 0.228 0.207 0.18 0.151 0.138 0.137 0.137
Noise 0.266 0.264 0.261 0.255 0.246 0.231 0.209 0.179 0.15 0.138 0.137 0.138
FGSM 0.265 0.263 0.258 0.251 0.239 0.221 0.197 0.167 0.144 0.137 0.137 0.138
PGD 0.265 0.262 0.257 0.248 0.235 0.216 0.19 0.161 0.141 0.137 0.138 0.138
AAE-Cox 0.269 0.269 0.269 0.268 0.265 0.259 0.25 0.231 0.183 0.138 0.137 0.137
SAWAR 0.14 0.139 0.138 0.138 0.137 0.137 0.137 0.137 0.137 0.137 0.137 0.137
Framingham DRAFT 0.831 0.825 0.811 0.783 0.724 0.618 0.459 0.289 0.174 0.124 0.114 0.11
Noise 0.836 0.834 0.83 0.82 0.79 0.709 0.543 0.331 0.185 0.126 0.115 0.11
FGSM 0.836 0.836 0.835 0.831 0.817 0.765 0.605 0.336 0.162 0.118 0.113 0.112
PGD 0.836 0.836 0.836 0.831 0.812 0.74 0.555 0.295 0.146 0.115 0.113 0.112
AAE-Cox 0.836 0.836 0.836 0.836 0.836 0.836 0.834 0.775 0.385 0.116 0.109 0.108
SAWAR 0.17 0.155 0.144 0.134 0.127 0.122 0.117 0.114 0.112 0.111 0.111 0.111
LeukSurv DRAFT 0.206 0.206 0.205 0.205 0.203 0.2 0.195 0.185 0.17 0.158 0.154 0.153
Noise 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.2 0.171 0.166 0.173
FGSM 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.203 0.187 0.161 0.158 0.159
PGD 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.202 0.183 0.16 0.157 0.158
AAE-Cox 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.198 0.162 0.154 0.152
SAWAR 0.194 0.19 0.184 0.178 0.171 0.165 0.16 0.155 0.151 0.148 0.147 0.146
TRACE DRAFT 0.62 0.611 0.593 0.558 0.504 0.432 0.347 0.265 0.204 0.172 0.165 0.162
Noise 0.627 0.626 0.623 0.611 0.579 0.518 0.426 0.317 0.225 0.176 0.166 0.162
FGSM 0.627 0.625 0.619 0.596 0.545 0.46 0.351 0.248 0.188 0.167 0.163 0.162
PGD 0.627 0.625 0.617 0.593 0.54 0.455 0.347 0.247 0.188 0.167 0.164 0.162
AAE-Cox 0.627 0.627 0.627 0.627 0.627 0.623 0.588 0.471 0.284 0.172 0.165 0.163
SAWAR 0.324 0.281 0.246 0.22 0.202 0.188 0.179 0.171 0.166 0.163 0.162 0.161
dataDIVAT1 DRAFT 0.702 0.689 0.663 0.614 0.527 0.405 0.285 0.21 0.177 0.17 0.172 0.177
Noise 0.72 0.72 0.719 0.718 0.713 0.689 0.599 0.4 0.245 0.197 0.193 0.195
FGSM 0.719 0.719 0.718 0.714 0.697 0.637 0.478 0.285 0.2 0.182 0.182 0.184
PGD 0.719 0.719 0.716 0.71 0.687 0.613 0.442 0.261 0.191 0.18 0.181 0.182
AAE-Cox 0.72 0.72 0.72 0.72 0.72 0.72 0.715 0.553 0.194 0.165 0.17 0.174
SAWAR 0.189 0.184 0.181 0.179 0.177 0.177 0.177 0.177 0.177 0.178 0.179 0.179
flchain DRAFT 0.816 0.816 0.816 0.816 0.816 0.816 0.814 0.797 0.46 0.093 0.069 0.054
Noise 0.816 0.816 0.816 0.816 0.816 0.816 0.814 0.77 0.227 0.088 0.067 0.057
FGSM 0.816 0.816 0.816 0.816 0.816 0.804 0.468 0.119 0.089 0.061 0.055 0.054
PGD 0.817 0.817 0.817 0.816 0.81 0.638 0.162 0.107 0.076 0.057 0.054 0.053
AAE-Cox nan nan nan nan nan nan nan nan 0.794 0.688 0.054 0.051
SAWAR 0.445 0.363 0.241 0.094 0.067 0.059 0.057 0.055 0.054 0.053 0.053 0.053
prostate DRAFT 0.517 0.517 0.516 0.513 0.508 0.495 0.466 0.408 0.322 0.234 0.202 0.181
Noise 0.518 0.518 0.519 0.519 0.52 0.52 0.519 0.508 0.457 0.321 0.267 0.249
FGSM 0.517 0.518 0.519 0.518 0.52 0.52 0.517 0.493 0.402 0.262 0.228 0.218
PGD 0.518 0.518 0.518 0.518 0.517 0.518 0.515 0.486 0.383 0.248 0.219 0.209
AAE-Cox 0.519 0.519 0.519 0.519 0.518 0.518 0.517 0.511 0.443 0.205 0.173 0.169
SAWAR 0.37 0.334 0.291 0.262 0.236 0.218 0.204 0.192 0.184 0.178 0.176 0.175
retinopathy DRAFT 0.728 0.722 0.71 0.687 0.647 0.579 0.48 0.364 0.266 0.204 0.188 0.179
Noise 0.73 0.725 0.714 0.693 0.652 0.579 0.472 0.353 0.256 0.199 0.184 0.177
FGSM 0.725 0.715 0.697 0.665 0.61 0.526 0.419 0.314 0.237 0.196 0.186 0.182
PGD 0.724 0.714 0.695 0.662 0.606 0.521 0.413 0.309 0.235 0.195 0.186 0.182
AAE-Cox 0.733 0.733 0.733 0.733 0.733 0.732 0.716 0.596 0.281 0.183 0.181 0.181
SAWAR 0.588 0.548 0.499 0.44 0.378 0.317 0.265 0.224 0.197 0.184 0.181 0.181
stagec DRAFT 0.556 0.549 0.544 0.543 0.539 0.517 0.468 0.404 0.332 0.274 0.256 0.245
Noise 0.559 0.553 0.547 0.548 0.551 0.541 0.501 0.437 0.357 0.282 0.258 0.244
FGSM 0.547 0.545 0.549 0.553 0.545 0.511 0.457 0.391 0.323 0.278 0.265 0.258
PGD 0.547 0.546 0.549 0.553 0.542 0.507 0.453 0.388 0.322 0.278 0.267 0.26
AAE-Cox 0.568 0.568 0.568 0.568 0.567 0.559 0.545 0.544 0.474 0.291 0.268 0.263
SAWAR 0.521 0.497 0.469 0.436 0.401 0.365 0.334 0.303 0.276 0.256 0.249 0.243
zinc DRAFT 0.847 0.847 0.846 0.843 0.831 0.787 0.641 0.376 0.184 0.119 0.109 0.107
Noise 0.847 0.847 0.847 0.845 0.838 0.795 0.627 0.351 0.179 0.128 0.118 0.113
FGSM 0.847 0.847 0.845 0.835 0.782 0.611 0.356 0.189 0.133 0.116 0.113 0.112
PGD 0.847 0.847 0.845 0.834 0.777 0.6 0.341 0.182 0.132 0.116 0.113 0.112
AAE-Cox 0.847 0.847 0.847 0.847 0.847 0.847 0.846 0.823 0.507 0.119 0.109 0.108
SAWAR 0.654 0.532 0.401 0.289 0.21 0.161 0.133 0.118 0.111 0.109 0.109 0.109